# Monte Carlo sampling with integrator snippets

### v1

@@ 4.1 Variance using folded mixture samples
4.2 Variance using unfolded mixture samples
@@
4.2.1 Relative efficiency for unfolded estimators
@@
For clarity we reproduce the very useful lemma [38, Lemma 6], correcting a couple of minor typos along the way.
@@ 4.2.2 Optimal weights
4.3 More on variance reduction and optimal flow
@@
We now show that in the examples considered in this paper our approach can be understood as implementing unbiased Riemann sum approximations of line integrals along contours. Adopting a continuous time approach can be justified as follows. For numerous integrators, the conditions of the following proposition are satisfied; this is the case for the leapfrog integrator of Hamilton’s equation e.g. [20, 38, Theorem 3.4] and [38, Appendix 3.1, Theorem 9] for detailed results.
@@
4.3.1 Hamiltonian contour decomposition
@@
Since $H^{-1}(s)$ is rectifiable we have that [9, Theorem 2.6.2] for $g\colon\mathsf{Z}\rightarrow\mathbb{R}$,
@@
where the limit always exists in the ${\rm L}^{2}(\mu)$ sense and constitutes Von Neumann’s mean ergodic theorem [41, 35], and the rightmost equality forms Boltzman’s conjecture. An interesting property in the present context is that $\mathbb{E}_{\bar{\mu}}\bigl{(}\bar{f}(Z)\bigr{)}=\mathbb{E}_{\mu}\bigl{(}f(Z)\bigr{)}$ for $f\in\mathrm{L}^{2}(\mu)$ and one could replicate the variance reduction properties developed earlier. Boltzman’s conjecture has long been disproved, as numerous Hamiltonians can be shown not to lead to ergodic systems, although some sub-classes do. However a weaker, or local, form or ergodicity can hold on sets of a partition of $\mathsf{Z}$
@@
where we have used a polar reparametrization $v=\rho\mathbf{e}$ $(\rho,\mathbf{e})\in\mathbb{R}_{+}\times\mathcal{S}_{d-1}$, and whether a marginal version of Boltzman’s conjecture holds for this limit. Example 25 indicates that this is not true in general but the cardinality of the partition of $\mathsf{X}$ may be reduced.
4.3.2 Advantage of averaging and control variate interpretation
@@
and we are interested in determining $t\mapsto\psi_{t}$ (or $F$) in order to minimize ${\rm var}_{\bar{\mu}}\left(\mathbb{E}_{\bar{\mu}}\big{(}\bar{f}(\mathsf{T},\check{Z})\mid\check{Z}\big{)}\right)$ and maximize improvement over simple Monte Carlo. We recall that the Jabobian determinant of the flow $t\mapsto\psi_{t}(z)$ is given by [19, p. 174108-5]
@@
4.3.3 Towards an optimal flow?
@@
5 MCMC with integrator snippets
We restrict this discussion to integrator snippet based algorithms, but more general Markov snippet algorithms could be considered.
Consider again the target distribution
@@ 5 MCMC with integrator snippets
where $\mu_{k}({\rm d}z):=\mu^{\psi^{-k}}({\rm d}z)$ for $k\in\llbracket 0,T\rrbracket$. Assume that we are in the context of Example 1, dropping $n$ for simplicity, the HMC algorithm using the integrator $\psi^{s}$ is as follows
@@
described algorithmically in Alg. 6. Indeed for any $(l,v^{\prime})\in\llbracket 0,T\rrbracket\times\mathsf{V}$, using identity (11),
@@
6 Discussion
We have shown how mappings used in various Monte Carlo schemes relying on numerical integrators of ODE can be implemented to fully exploit all computations to design robust and efficient sampling algorithms. Numerous questions remain open, including the tradeoff between $N$ and $T$. A precise analysis of this question is made particularly difficult by the fact that integration along snippets is straightforwardly parallelizable, while resampling does not lend itself to straightforward parallelisation.
We leave all these questions for future investigations.
Acknowledgements
The authors would like to thank Carl Dettman for very useful discussions on Boltzman’s conjecture. Research of CA and MCE supported by EPSRC grant ‘CoSInES (COmputational Statistical INference for Engineering and Security)’ (EP/R034710/1), and EPSRC grant Bayes4Health, ‘New Approaches to Bayesian Data Science: Tackling Challenges from the Health Sciences’ (EP/R018561/1). Research of CZ was supported by a CSC Scholarship.
Appendix A Notation and definitions
  • For $A\subset\mathsf{E}$ we let $A^{\complement}$ be its complement.
  • For a Markov kernel $P(x,{\rm d}y)$ on $\mathsf{E}\times\mathscr{E}$, we write
    $\int f(x)\delta_{x}({\rm d}y)=f(x)$
    $\sum_{i=1}^{N}\frac{w_{i}}{\sum_{j=1}^{N}w_{j}}f(z_{i})\approx\mu(f)\,,$
    in either in the $L^{p}$ sense for some $p\geq 1$.
  • for $f\colon\mathbb{R}^{m}\rightarrow\mathbb{R}^{n}$ we let
    • $\nabla\otimes f$ be the transpose of the Jacobian
    • for $n=1$ we let $\nabla f=(\nabla\otimes f)^{\top}$ be the gradient,
    • $\nabla\cdot f$ be the divergence.
Appendix B Radon-Nikodym derivative
Definition 28 (Pushforward).
$\mu^{\psi}(A):=\mu(\psi^{-1}(A)),\qquad A\in\mathscr{Z}^{\prime},$
Definition 29 (Dominating and equivalent measures).
  1. 1.
  2. 2.
We will need the notion of Radon-Nikodym derivative [7, Theorems 32.2 & 16.11]:
Theorem 30 (Radon–Nikodym).
$\int_{A}f(z)\mu({\rm d}z)=\nu(A),\qquad A\in\mathscr{E}.$
The following establishes the expression of an expectation with respect to the pushforward $\mu^{\psi}$ in terms of expectations with respect to $\mu$ [7, Theorem 16.13].
Theorem 31 (Change of variables).
Lemma 32.
  1. 1.
  2. 2.
    for $\upsilon$-almost all $z\in\mathsf{Z}$,
  3. 3.
    we have
    in which case for $\upsilon$-almost all $z\in\mathsf{Z}$
    and therefore
Proof.
$=\int f\circ\psi^{-1}(z)\mu({\rm d}z)$
The third statement is given as [7, Problem 32.6.], which we solve in Lemma 34. ∎
Corollary 33.
Lemma 34 (Billingsley, Problem 32.6.).
Proof.
$=$ $0\,,$
and we conclude. ∎

### v2

@@ H.1 Variance using folded mixture samples
H.2 Variance using unfolded mixture samples
@@
H.2.1 Relative efficiency for unfolded estimators
@@
For clarity we reproduce the very useful lemma [61, Lemma 6], correcting a couple of minor typos along the way.
@@ H.2.2 Optimal weights
H.3 More on variance reduction and optimal flow
@@
We now show that in the examples considered in this paper our approach can be understood as implementing unbiased Riemann sum approximations of line integrals along contours. Adopting a continuous time approach can be justified as follows. For numerous integrators, the conditions of the following proposition are satisfied; this is the case for the leapfrog integrator of Hamilton’s equation e.g. [35, 61, Theorem 3.4] and [61, Appendix 3.1, Theorem 9] for detailed results.
@@
H.3.1 Hamiltonian contour decomposition
@@
Since $H^{-1}(s)$ is rectifiable we have that [14, Theorem 2.6.2] for $g\colon\mathsf{Z}\rightarrow\mathbb{R}$,
@@
where the limit always exists in the ${\rm L}^{2}(\mu)$ sense and constitutes Von Neumann’s mean ergodic theorem [65, 58], and the rightmost equality forms Boltzmann’s conjecture. An interesting property in the present context is that $\mathbb{E}_{\bar{\mu}}\bigl{(}\bar{f}(Z)\bigr{)}=\mathbb{E}_{\mu}\bigl{(}f(Z)\bigr{)}$ for $f\in\mathrm{L}^{2}(\mu)$ and one could replicate the variance reduction properties developed earlier. Boltzmann’s conjecture has long been disproved, as numerous Hamiltonians can be shown not to lead to ergodic systems, although some sub-classes do. However a weaker, or local, form or ergodicity can hold on sets of a partition of $\mathsf{Z}$
@@
where we have used a polar reparametrization $v=\rho\mathbf{e}$ $(\rho,\mathbf{e})\in\mathbb{R}_{+}\times\mathcal{S}_{d-1}$, and whether a marginal version of Boltzmann’s conjecture holds for this limit. Example 33 indicates that this is not true in general but the cardinality of the partition of $\mathsf{X}$ may be reduced.
H.3.2 Advantage of averaging and control variate interpretation
@@
and we are interested in determining $t\mapsto\psi_{t}$ (or $F$) in order to minimize ${\rm var}_{\bar{\mu}}\left(\mathbb{E}_{\bar{\mu}}\big{(}\bar{f}(\mathsf{T},\check{Z})\mid\check{Z}\big{)}\right)$ and maximize improvement over simple Monte Carlo. We recall that the Jabobian determinant of the flow $t\mapsto\psi_{t}(z)$ is given by [31, p. 174108-5]
@@
H.3.3 Towards an optimal flow?
@@
Appendix I MCMC with integrator snippets
We restrict this discussion to integrator snippet based algorithms, but more general Markov snippet algorithms could be considered. Consider again the target distribution
@@ Appendix I MCMC with integrator snippets
where $\mu_{k}({\rm d}z):=\mu^{\psi^{-k}}({\rm d}z)$ for $k\in\llbracket 0,T\rrbracket$.
I.1 Windows of states
Assume that we are in the context of Example 1, dropping $n$ for simplicity, the HMC algorithm using the integrator $\psi^{s}$ is as follows
@@ I.1 Windows of states
described algorithmically in Alg. 8. Indeed for any $(l,v^{\prime})\in\llbracket 0,T\rrbracket\times\mathsf{V}$, using identity (10),
@@
I.2 Multinomial HMC
Multinomial HMC is another related MCMC update suggested in [9] which can be rephrased in terms of our framework as follows. For $(z,A)\in\mathsf{Z}\times\mathscr{Z}$ define the kernel
$P(z,A)=\frac{1}{T+1}\sum_{k,l\in\llbracket 0,T\rrbracket}\bar{\mu}\big{(}k\mid\psi^{-l}(z)\big{)}\mathbf{1}\big{\{}\psi^{k-l}(z)\in A\big{\}}\,.$
Introducing velocity refreshment before returning $\psi^{k-l}(z)$, can be reframed in terms $\bar{M}\colon\mathsf{Z}\times\mathscr{Z}$ as defined in (10) leading the the Markov kernel
$P^{\prime}(z,A)=\frac{1}{T+1}\sum_{l\in\llbracket 0,T\rrbracket}\bar{M}\big{(}\psi^{-l}(z),A\big{)}\,.$
We focus below on the version with no refreshment, $P$. Establishing $\mu$-invariance is direct since with $z\sim\mu$ and $l\sim\mathcal{U}\big{(}\llbracket 0,T\rrbracket\big{)}$ then $\psi^{-l}(z)\sim\bar{\mu}$ and with $k\sim\bar{\mu}\big{(}\cdot\mid\psi^{-l}(z)\big{)}$ we know that $\psi^{k-l}(z)\sim\mu$ from (10). More formally
$\mu P(A)$ $=\frac{1}{T+1}\sum_{k\in\llbracket 0,T\rrbracket}\int\mu({\rm d}z)\sum_{l\in\llbracket 0,T\rrbracket}\bar{\mu}\big{(}k\mid\psi^{-l}(z)\big{)}\mathbf{1}\big{\{}\psi^{k}\circ\psi^{-l}(z)\in A\big{\}}$
$=\sum_{k\in\llbracket 0,T\rrbracket}\int\bar{\mu}({\rm d}z)\bar{\mu}\big{(}k\mid z\big{)}\mathbf{1}\big{\{}\psi^{k}(z)\in A\big{\}}$
$=\mu(A)\,.$
This Markov kernel is in fact reversibility, that is for $f,g\colon\mathsf{Z}\rightarrow\mathbb{R}$
$\sum_{k,l\in\llbracket 0,T\rrbracket}\int f(z)\mu({\rm d}z)\bar{\mu}\big{(}k\mid\psi^{-l}(z)\big{)}g\circ\psi^{k-l}(z)$
$=\frac{1}{T+1}\sum_{k,l\in\llbracket 0,T\rrbracket}\int$ $f\circ\psi^{l}(z)\mu^{\psi^{-l}}({\rm d}z)\frac{{\rm d}\mu^{\psi^{-k}}}{{\rm d}\bar{\mu}}\big{(}z\big{)}g\circ\psi^{k}(z)$
$=\frac{1}{T+1}\sum_{k,l\in\llbracket 0,T\rrbracket}\int$ $f\circ\psi^{l}(z)\frac{{\rm d}\mu^{\psi^{-l}}}{{\rm d}\bar{\mu}}(z)\mu^{\psi^{-k}}\big{(}{\rm d}z\big{)}g\circ\psi^{k}(z)$
$=\frac{1}{T+1}\sum_{k,l\in\llbracket 0,T\rrbracket}\int$ $f\circ\psi^{l-k}(z)\frac{{\rm d}\mu^{\psi^{-l}}}{{\rm d}\bar{\mu}}\circ\psi^{-k}(z)\mu\big{(}{\rm d}z\big{)}g(z)$
$=\sum_{k,l\in\llbracket 0,T\rrbracket}\int$ $f\circ\psi^{l-k}(z)\mu({\rm d}z)\bar{\mu}\big{(}l\mid\psi^{-k}(z)\big{)}g(z)$
A standard SMC implementation relying on this kernel and corresponding near optimal backward kernel would lead to an algorithm different from ours. Indeed in this scenario, for $n\in\llbracket P\rrbracket$, for particles $\{\dot{z}_{n}^{(i)},i\in\llbracket N\rrbracket\}$ the weights are of the standard form $\{{\rm d}\mu_{n+1}/{\rm d}\mu_{n}(\dot{z}_{n+1}^{(i)}),i\in\llbracket N\rrbracket\}$, which should be contrasted with the weights in Alg. (6) $\big{\{}{\rm d}\bar{\mu}_{n+1}/{\rm d}\mu_{n}\big{(}z_{n}^{(i)}\big{)},i\in\llbracket N\rrbracket\big{\}}$ where resampling “looks ahead” and uses the information provided by the snippets stemming from $\big{\{}z_{n}^{(i)},i\in\llbracket N\rrbracket\big{\}}$. Note that there is no indexing mistake here and that the additional dots have been added for coherence: in Alg. (6) the output of iteration $n+1$ are $\{\check{z}_{n+1}^{(i)},i\in\llbracket N\rrbracket\}$ and $\big{\{}z_{n}^{(i)},i\in\llbracket N\rrbracket\big{\}}$ intermediate states to which resampling is applied. One expects robustness to arise from this difference.